OpenStudy (anonymous):

A projectile is fired at time t=0.0s, from point 0 at the edge of a cliff, with initial velocity components of vox=80m/s and voy=200m/s. The projectile rises, then falls into the sea at point P. The time of flight of the projectile is 50.0s WHEN THE Y COMPONENT OF VELOCITY EQUALS 160m/s WHAT IS THE X COMPONENT A. 280m B. 290m C. 310m D. 340m

OpenStudy (shamim):

Can u draw it ?

OpenStudy (jiteshmeghwal9):

Let us find the time when the velocity of Y_component becomes 160 m/s. \[v_y=u_y-g t\]\[160=200-10t\]\[t=4 \sec\]As we know x-component of velocity will remain same throughout the projectile motion. Therefore,\[\Large{X=u_xt}\]u_x=80 m/s t=4 sec. solve it

OpenStudy (anonymous):

Thanks guys, the equations helped